OpenStudy (anonymous):

prove the following trigonometric identities. 2/(rt(3)cos(x)+sin(x))=sec(pi/6-x) tan(x/2)=sin(x)/(1+cos(x))

OpenStudy (anonymous):

@dinamix

OpenStudy (anonymous):

\[\frac{ 2 }{\sqrt{3} \cos(x)+\sin(x)}=\sec(\pi/6-x) \tan(x/2)=\frac{ \sin (x) }{1+\cos (x) } \] is that your question?

OpenStudy (anonymous):

no they are two separate equations. 2/(rt(3)cos(x)+sin(x))=sec(pi/6-x) tan(x/2)=sin(x)/(1+cos(x))

OpenStudy (anonymous):

oK

OpenStudy (anonymous):

I will do only one, which one do you like?

OpenStudy (anonymous):

do the top one

OpenStudy (anonymous):

you can do either one if you want

OpenStudy (irishboy123):

for the second \[tan\frac{x}{2}=\frac{sinx}{1+cosx}\] try this \[tan\frac{x}{2}=\frac{2sin\frac{x}{2}cos \frac{x}{2}}{1+(1-2sin^2 \frac{x}{2})}\]

OpenStudy (welshfella):

sec (pi/6 - x) = 1 / (sqrt3/2 cos x + sin pi/6 sin x) = 2 / (sqrt 3 cos + sin x)

OpenStudy (welshfella):

* I missed out the second step which is = 1 / (cos pi/6 cos x + sin pi/6 sin x)

OpenStudy (anonymous):

\[\frac{ 2 }{\sqrt{3} \cos(x)+\sin(x)}=\sec(\pi/6-x)\] \[\frac{ 2 }{\sqrt{3} \cos(x)+\sin(x)}=\frac{ 1 }{ \cos{(\pi/6-x)} }\] using:cos(u-v)=cos u cos v+sin u sin v \[\frac{ 2 }{\sqrt{3} \cos(x)+\sin(x)}=\frac{ 1 }{ \cos{(\pi/6) \times \cos(x)}+\sin{(\pi/6) \times \sin(x)} }\]

OpenStudy (anonymous):

i got that far ASAAD123 put i didn't know what to do after that

OpenStudy (welshfella):

the denominator = cos (pi/6 - x)

OpenStudy (welshfella):

and 1 / cos = sec

OpenStudy (anonymous):

wouldn't you also change sin(pi/6) to 1/2

OpenStudy (welshfella):

no its the sin of a compound angle (pi/6 - x) not sin pi/6

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so how do you get the sides to equal

OpenStudy (welshfella):

i had to use sin pi/6 = 1/2 because i started with RHS and converted to LHS whereas ASAAD did the reverse

OpenStudy (welshfella):

with these identities you choose one side and try to convert it to the other. This proves the identity.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i think i understand it now

OpenStudy (welshfella):

so you can use asaad's or mine. Either would do.

OpenStudy (welshfella):

There is only more more step to prove IrishBoy's solution

OpenStudy (anonymous):

i already got irish boy's solution